Given 5x+2y-6=0 write equation of a line that is perpendicular going through point (0,-3).

Found 2 solutions by fractalier, stanbon:
Answer by fractalier(6550)

Find the slope of this line first.
5x + 2y - 6 = 0
2y = -5x + 6
y = (-5/2)x + 3
This time you need a slope of 2/5 (it's the negative reciprocal)...
So using the point and y = mx + b, we get
-3 = (2/5)(0) + b
and b = -3
so your line is
y = (2/5)x - 3

Answer by stanbon(75887)

Given 5x+2y-6=0 write equation of a line that is perpendicular going through point (0,-3).
-------------
Slope of the given line::
y = (-5/2)x + 3
slope = -5/2
---------
Slope of the perpendicular line:: 2/5
-----
Form of the answer::
y = mx + b
y = (2/5)x -3
-----------------
Cheers,
Stan H.
-------------